OpenStudy (anonymous):

A toaster was originally $30. It is on sale for 25% off. What is the discounted price?

OpenStudy (anonymous):

Do you mean what is the price of the toaster after the discount?

OpenStudy (anonymous):

Thats what the question asks me

OpenStudy (anonymous):

Ok well the answer is going to be 30 - X where X is what you save x= 30*.25

OpenStudy (anonymous):

30- (30*0.25)=?

OpenStudy (anonymous):

Can you solve it now?

OpenStudy (anonymous):

25/100% x$30 =$7.5 So, $30-$7.5=$22.5 $22.5 is the price after discount.

OpenStudy (anonymous):

@jishalini Please don't give answers like that

OpenStudy (anonymous):

What's wrong with my answer?

OpenStudy (anonymous):

You gave it away

OpenStudy (anonymous):

You sniped me.

OpenStudy (anonymous):

Oh.. I'm sorry. Didn't mean to do so. Sorry again.

OpenStudy (anonymous):

If someone if helping another person you can help as well....but don't give out the final answer.

OpenStudy (anonymous):

Sometimes it confuses the asker that's why